given a 6 percent interest rate, compute the present value of payments made in years 1,2,3 and 4 of $1,100, $1400, $1400, and %1500 respectively
Answer #1

Answer:-

The interest rate (r) = 6%

The present value = Amount / (1+r)n

The first year payment =$ 1100
The present value = $ 1100 / 1.06 = $ 1037.74

The second year payment =$ 1400
The present value = $ 1400 / 1.062 = $ 1245.99

The third year payment =$ 1400
The present value = $ 1400 / 1.063 = $ 1175.47

The fourth year payment =$ 1500
The present value = $ 1500 / 1.064 = $ 1188.59